ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

13. Combinatorics

1991 MS Classification: 53 Alli, Vahid, University of Imam Hossein, Tehran, Iran Totally Geodesic Subspace of a Finsler Space This is well known that a totally geodesic (or totally umblical) subspace M_{m} of a Riemannian space M_{n} is of constant curvature. In this paper an effort has been made to study the same result in Finsler Space and to investigate the properties of hv and v-curvature tensors of M_{m}. In the last the totally geodesic of first, second and third kinds have been discussed.